VIBRATIONAL OVERTONE SPECTRA OF C_{2}H_{6} AND C_{2}H_{4} IN CRYOGENIC LIQUIDS

\underline{MARIA H. DIEZ-Y-RIEGA} and CARLOS E. MANZANARES, Department of Chemistry and Biochemistry, Baylor University, Waco, Texas, 76798.

Vibrational overtone spectra of C_{2}H_{6} and C_{2}H_{4} in cryogenic solutions were recorded between 5000 and 14000 cm^{−1}. Spectral regions for the first four overtones were measured using a Fourier transform spectrophotometer. The fifth overtone (Δν = 6) spectra between 15,000 and 16,000 cm^{−1} were recorded with a double beam (pump-probe) thermal lens technique using concentrations as low as 10^{−3} mole fraction. The peak frequency shift (Δω) from gas phase to solution is explained by the change in harmonic frequency and anharmonicity in solution with respect to the gas phase values. The bandwidth (Δω_{1/2}) of the (Δν = 6) C-H absorption bands in solution can be explained in terms of collisions with the solvent molecules.